Please graph the following functions in a separate piece of paper. I y = 3 tan (2x x- a) 2 TT y = - tan ( x + b) 4
Answer #1

Solution! y = 3 tan (2x - TL Here Amplitude: None period :I phase Shift 퓨 : IT Vertical Shift - The graph is w I Go to -> X -b y = - tan (x+1) Here period - phase shift = Amplitude - None y - X 31. NI 1-311 - 플 Thanks! please like